OpenStudy (anonymous):

Please walk me through this; I am a bit lost with this one! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

she starts with $17

jimthompson5910 (jim_thompson5910):

after year 1, how much interest does she earn?

OpenStudy (anonymous):

A= 19.244 ?

jimthompson5910 (jim_thompson5910):

it says "simple interest" so we use I = P*r*t

OpenStudy (anonymous):

2.24

jimthompson5910 (jim_thompson5910):

I = P*r*t I = 17*0.011*1 I = 0.187 I = 0.19 so the interest after 1 yr is 19 cents

jimthompson5910 (jim_thompson5910):

how about after 2 years?

OpenStudy (anonymous):

4.488 =4.49

jimthompson5910 (jim_thompson5910):

idk how you're getting that

OpenStudy (anonymous):

17*0.011*24= 4.488

jimthompson5910 (jim_thompson5910):

t is time in years, not months

jimthompson5910 (jim_thompson5910):

t = 2 for 2 yrs (you don't use 24)

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

0.374

jimthompson5910 (jim_thompson5910):

so she'll have 17+0.37 = 17.37 in the account at year 2

jimthompson5910 (jim_thompson5910):

the point (2, 17.37) represents this

OpenStudy (anonymous):

Are we testing more options out or we've concluded it's B ?

jimthompson5910 (jim_thompson5910):

you can test others and I recommend you do

jimthompson5910 (jim_thompson5910):

other points I mean, like what is the amount in the account at year 6?

OpenStudy (anonymous):

I wouldn't prefer/like to, I'm just wondering if B would be the final answer?

jimthompson5910 (jim_thompson5910):

well it helps with practice

OpenStudy (anonymous):

Yes, progress.. Thank you soooooo much for everything!! Thank you for ****A L L**** your help!!!! ~OVER-Abundant blessings and good vibes your way!!!! ^^.^^